# Measure gradients, not activations! Enhancing neuronal activity in deep reinforcement learning

### v1

@@ Abstract
Deep reinforcement learning (RL) agents frequently suffer from neuronal activity loss, which impairs their ability to adapt to new data and learn continually. A common method to quantify and address this issue is the $\tau$-dormant neuron ratio, which uses activation statistics to measure the expressive ability of neurons. While effective for simple MLP-based agents, this approach loses statistical power in more complex architectures. To address this, we argue that in advanced RL agents, maintaining a neuron’s learning capacity, its ability to adapt via gradient updates, is more critical than preserving its expressive ability. Based on this insight, we shift the statistical objective from activations to gradients, and introduce GraMa (Gradient Magnitude Neural Activity Metric), a lightweight, architecture-agnostic metric for quantifying neuron-level learning capacity. We show that GraMa effectively reveals persistent neuron inactivity across diverse architectures, including residual networks, diffusion models, and agents with varied activation functions. Moreover, resetting neurons guided by GraMa (ReGraMa) consistently improves learning performance across multiple deep RL algorithms and benchmarks, such as MuJoCo and the DeepMind Control Suite.
@@ 1 Introduction
Deep reinforcement learning (Deep RL) has achieved remarkable success across a variety of domains, including robotics (Liu et al., 2021), drone control (Song et al., 2021), and game playing (Berner et al., 2019, Schwarzer et al., 2023). These advancements have been driven by the expressive power and adaptive learning ability of neural networks which effectively approximate and optimize value functions and/or policies (Sokar et al., 2023). However, recent studies have uncovered a critical and often underexplored challenge: as training progresses, subsets of neurons in these networks often experience a progressive loss of activity and become dormant (Sokar et al., 2023, Ma et al., 2024, Qin et al., 2024). This phenomenon reduces the learning capacity of the network, comprising its ability to adapt to non-stationary data distributions (Nikishin et al., 2022), which in turn hinders their ability to acquire new knowledge and adapt to evolving environments (Abbas et al., 2023). Despite its importance, quantifying and mitigating neuronal activity remains challenging due to its complex underlying mechanisms (Lyle et al., 2023, Nauman et al., 2024a, Lyle et al., 2024, Obando Ceron et al., 2023).
@@
To address this problem, a primary principle has been to restore a network’s learning ability by reactivating or resetting inactive neurons. These approaches span multiple granularities: model-level and layer-level resets that reinitialize specific layers, while straightforward, often lead to catastrophic forgetting (Nikishin et al., 2023). On the other hand, neuron-dependent resets (e.g., ReDo (Sokar et al., 2023)) target specific underperforming neurons (Xu et al., 2023) and provide a finer-grained approach by selectively reinitializing a subset of neurons identified as dormant, which mitigates forgetting and maintains computational efficiency.
@@
The effectiveness of neuron-dependent resets hinges critically on having a reliable criterion to identify which neurons require initialization. Existing methods primarily rely on activation-based metrics, such as the $\tau$ dormant neuron metric (Sokar et al., 2023), which measures neuronal inactivity based on activation values (i.e., the output of activation functions). It has demonstrated utility in standard architectures, and has been used to guide targeted neuron resets to restore activity in simple settings such as serial MLPs with ReLU activations (Agarap, 2018), providing a simple means of maintaining learning capability without relying on auxiliary networks (Nikishin et al., 2023) or models (Lee et al., 2024). By identifying neurons with weak or no activation and selectively reinitializing them, the activation-based dormancy metrics become a widely adopted tool for restoring learning capabilities and preventing performance plateau in standard deep RL settings (Farias and Jozefiak, 2025, Liu et al., 2025, Juliani and Ash, 2024).
@@
Figure 1: Dormant neuron metric struggles in advanced vision RL BRO-net agent (Nauman et al., 2024b). Neuron resetting based on the dormant neuron index (ReDo) cannot restore the learning capacity of the agent, which limits its effectiveness. The curves record normalized score across 3 image-input tasks (15 runs per method), i.e., Dog Stand, Dog Walk, Dog Run.
@@
However, deep RL architectures have rapidly evolved beyond these simple architectures, integrating advanced components such as residual connections (Nauman et al., 2024b, Lee et al., 2025a), mixture of experts (MoEs) (Obando-Ceron et al., 2024, Sokar et al., 2024, Willi et al., ), and diffusion-based models (Ren et al., 2025, Wang et al., 2024) to improve scalability and performance for tackling complex tasks involving continuous control and high-dimensional observations. Additionally, they also imperceptibly alter the statistical properties of neuron activations, creating a significant mismatch with traditional activity metrics. Through our extensive analysis, we find that resetting based on $\tau$-dormant neuron ratio (ReDo) reduces effectiveness in these newer architectures. As shown in Figure 1, ReDo struggles to identify inactive neurons that have lost their learning capacity in the more advanced BRO-net architecture, which features residual paths and layer normalization, ultimately failing to achieve the intended enhancement. This mismatch arises because activation-based metrics focus solely on a neuron’s current output, evaluating its expressive capacity (how strongly it activates), while neglecting its learning capacity (how effectively it can adapt to new data distributions). Consequently, it becomes particularly problematic as deep RL architectures evolve beyond simple feedforward networks to incorporate advanced structural elements, diverse activation functions, and sophisticated normalization techniques. In these modern architectures, a neuron’s activation magnitude often fails to accurately reflect its true learning potential.
@@
To address this mismatch, we propose a fundamental shift in perspective: from evaluating neurons based on their outputs to evaluating their learning potential by leveraging gradient magnitude. While activation values only capture what a neuron currently expresses, gradients tend to measure a neuron’s capacity in response to the given situation that directly drives parameter updates. This makes it a general and natural proxy for neuronal health across architectural variations. Building upon this insight, we introduce GraMa (Gradient Magnitude based Neuronal Activity Metric), a robust and lightweight framework for quantifying neuronal activity via the gradient magnitudes. GraMa maintains validity across diverse architectural patterns, making it well-suited for modern deep RL agents. In addition, GraMa imposes negligible computational and memory overhead by utilizing information already present in the optimization pipeline, which is a critical consideration for resource-intensive RL training. Leveraging GraMa’s efficiency, we develop a targeted neuron reset mechanism, (ReGraMa), that selectively reinitializes inactive neurons that have lost their learning capacity during training. This mechanism demonstrates robust efficiency across various architectures, including the SAC variant (Nadimpalli et al., 2025), the residual BRO-net (Nauman et al., 2024b), and the diffusion-based policy DACER (Wang et al., 2024). Our contributions are summarized as follows:
  • We show that the widely-adopted activation-based neuronal health measurements lose statistical power in complex architectures and provide a qualitative analysis of the underlying causes.
  • We reframe neuronal health evaluation through Grama, a gradient-based metric that quantifies learning potential independently of architectural complexity, and demostrate that neuronal learning capacity degradation affects even state-of-the-art network developments.
  • We develop (ReGraMa), an efficient neuron resetting mechanism guided by GraMa, which effectively restores neuronal activity across a wide range of network architectures.
  • We conduct extensive experiments on MuJoCo (Brockman et al., 2016), DeepMind Control Suite (Tassa et al., 2018), showing that GraMa-guided resetting improves performance and learning stability across diverse architectures.
@@ 2 Background
A reinforcement learning (RL) problem is typically formalized as a Markov Decision Process (MDP), defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathcal{P}$ the transition probability function $\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$, $\mathcal{R}$ the reward function $\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, and $\gamma\in[0,1)$ the discount factor. The state-action value function under policy $\pi$ is given by:$Q^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right].$ The objective is to find a policy $\pi$ that maximizes the expected return $Q^{\pi}(s,a)$ for each state $s$. In actor-critic-based deep RL, the Q-function is approximated by a neural network $Q_{\theta}$ with parameters $\theta$. During training, the agent interacts with the environment and stores trajectories in a replay buffer $D$. Mini-batches sampled from $D$ are used to update $Q_{\theta}$ by minimizing the temporal difference loss:$\mathcal{L}_{Q}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim D}\left[(Q_{\theta}(s,a)-Q^{\mathcal{T}}(s,a))^{2}\right],$ where the target is given by $Q^{\mathcal{T}}(s,a)=\mathcal{R}(s,a)+\gamma Q_{\bar{\theta}}(s^{\prime},\pi_{\bar{\phi}}(s^{\prime}))$, and $\bar{\theta},\bar{\phi}$ denote the parameters of target critic and actor networks, respectively.
@@
Recent studies have identified that the dynamic and non-stationary nature of RL objectives can cause neurons to permanently lose their activity (Lu et al., 2018), impairing the network’s ability to fit new data and thereby limiting learning progress (Nikishin et al., 2022, Ceron et al., 2024). This phenomenon, referred to as the dormant neuron phenomenon, has been quantitatively characterized using the $\tau-$dormant neuron ratio (Sokar et al., 2023), and serves as a core metric for assessing neuron-level plasticity (Xu et al., 2023, Qin et al., 2024, Liu et al., 2025). Specifically, a neuron $i$ in layer $\ell$ is considered $\tau$-dormant if its normalized activation (see Equation 1) falls below a threshold $\tau$, where $H^{\ell}$ is the number of neurons in layer $\ell$, and $D$ denotes the data distribution. $h_{i}^{\ell}(x)$ represents the activation value of neuron $i$ given input $x$.
@@ 3 Gradient Magnitude based Neuronal Activity Metric (GraMa)
In this section, we qualitatively investigate ReDo’s limitations from the perspective of architectural composition subsection 3.1. We then introduce GraMa (subsection 3.2), a novel neuronal activity metric that redefines the statistical objective from activation values to gradient magnitudes. Next, we perform an in-depth analysis of the characteristics of low learning capacity neurons using GraMa, and provide both theoretical and empirical analyses of the performance similarity between GraMa and ReDo on simple architectures. Finally, we analyze the advantages of GraMa in complex architectures.
@@ 3.1 Misalignment of activation-based neuron activity detection in modern architectures
The widely adopted activation-based neuron activity measures (Sokar et al., 2023) operate on the assumption that a neuron’s activation magnitude directly corresponds to its contribution to learning. Motivated by the development of more complex network structures in language and vision domains, recent trends in deep RL have moved beyond simple serial MLP architectures with ReLU activations towards more sophisticated network designs for scaling. However, this assumption becomes increasingly fragile in modern complex architectures which are rapidly becoming standard practice. We reveal a misalignment between activation values and actual learning potential, leading to inefficient neuron identification.
@@
To systematically analyze this misalignment, we evaluate BRO-net (Nauman et al., 2024b), a recently proposed architecture, on the challenging Dog Walk task (Tassa et al., 2018). We categorized neurons into four quadrants based on their expressive capacity (measured by activation magnitude) and learning capacity (measured by gradient magnitude), which we introduce in more detail in subsection 3.2. Neurons are ranked according to these two criteria, and we select the top 25% of neurons in each ranking, so as to provide an understanding of the types of neurons reset by each method. The results in Figure 3 demonstrate that ReDo cannot accurately identify neurons that have genuinely lost their learning capacity (red) based on their activation values. More importantly, ReDo mistakenly resets a significant number of neurons with high learning capacity that are merely less expressive at the moment (dark blue).
@@
In addition to the primary factor of statistical objectives, we identify three key architectural features that may undermine the reliability of activation-based neuron dormancy detection. Figure 2 provides an intuitive illustration of each. (i) Multi-branch network structures. In modern architectures like ResNets (He et al., 2015), information flows through multiple branches before being fused and passing through activation functions. This architectural pattern introduces a critical problem: the final activation values observed after branch fusion no longer accurately reflect the individual contributions of neurons in each branch. Our experiments with Resnet-SAC (Shah and Kumar, 2021) shown in Figure 4 (a) confirm this effect, where the results suggest that activation-based methods lose their ability to identify dormant neurons and limits the performance of the agent. (ii) Non-ReLU activation function. The interpretability of activation values is highly dependent on the specific activation function used. While ReLU creates a clear distinction between “dead” (output=00) and “active” neurons, this clarity breaks down when considering alternative activation functions. For example, with Leaky ReLU, neurons may remain expressive and contribute to learning even when their pre-activation values are negative. As a result, solely relying on activation values to measure neuron dormancy becomes unreliable and may lead to misidentification. Our experiments replacing ReLU with Leaky ReLU in SAC (Figure 4 (b)) demonstrate this problem, where the activation-based method (ReDo) shows considerable reset activity during early learning with lower performance, indicating that it struggles to establish meaningful dormancy criteria. (iii) Normalization layers. Normalization techniques modify the distribution of neuron outputs before they pass through activation functions. This process adjusts outliers and rescales values across entire layers, causing post-normalization values to lose a direct correspondence with individual neuron functionality. Thus, activation values measured after normalization may no longer accurately reflect a neuron’s learning capacity. Our experiments incorporating layer normalization into SAC, Figure 4 (c) reveal that the activation-based method struggles to maintain consistent neuron assessment, whereas our method consistently identifies relevant neurons throughout the learning process.
Figure 4: Empirical validation corresponding to the three cases of Figure 2. Top row records the proportion of reset neurons. Bottom row shows the performance. Each curve represents the average over 3 seeds (Dog Walk).

### v2

@@ Abstract
Deep reinforcement learning (RL) agents frequently suffer from neuronal activity loss, which impairs their ability to adapt to new data and learn continually. A common method to quantify and address this issue is the $\tau$-dormant neuron ratio, which uses activation statistics to measure the expressive ability of neurons. While effective for simple MLP-based agents, this approach loses statistical power in more complex architectures. To address this, we argue that in advanced RL agents, maintaining a neuron’s learning capacity, its ability to adapt via gradient updates, is more critical than preserving its expressive ability. Based on this insight, we shift the statistical objective from activations to gradients, and introduce GraMa (Gradient Magnitude Neural Activity Metric), a lightweight, architecture-agnostic metric for quantifying neuron-level learning capacity. We show that GraMa effectively reveals persistent neuron inactivity across diverse architectures, including residual networks, diffusion models, and agents with varied activation functions. Moreover, resetting neurons guided by GraMa (ReGraMa) consistently improves learning performance across multiple deep RL algorithms and benchmarks, such as MuJoCo and the DeepMind Control Suite. We make our code available [^1] Code: https://github.com/torressliu/grad-based-plasticity-metrics.
@@ 1 Introduction
Deep reinforcement learning (Deep RL) has achieved remarkable success across a variety of domains, including robotics (Liu2021DeepRL), foundation model fine-tuning (Shao2024DeepSeekMathPT; Liu2025PartIT; Yu2025DAPOAO; Li2025AttentionIL; Lu2025VLARLTM; Liu2025AsymmetricPP), and game playing (Berner2019Dota2W; pmlr-v202-schwarzer23a). These advancements have been driven by the expressive power and adaptive learning ability of neural networks which effectively approximate and optimize value functions and/or policies (sokar2023dormant). However, recent studies have uncovered a critical and often underexplored challenge: as training progresses, subsets of neurons in these networks often experience a progressive loss of activity and become dormant (sokar2023dormant; ma2024revisiting; qin2024the). This phenomenon reduces the learning capacity of the network, comprising its ability to adapt to non-stationary data distributions (Nikishin2022ThePB), which in turn hinders their ability to acquire new knowledge and adapt to evolving environments (Abbas2023LossOP). Despite its importance, quantifying and mitigating neuronal activity remains challenging due to its complex underlying mechanisms (understand; obando2023small; overfit; Lyle2024DisentanglingTC; liu2025the).
@@
To address this problem, a primary principle has been to restore a network’s learning ability by reactivating or resetting inactive neurons. These approaches span multiple granularities: model-level and layer-level resets that reinitialize specific layers, while straightforward, often lead to catastrophic forgetting (nikishin2023deep). On the other hand, neuron-dependent resets (e.g., ReDo (sokar2023dormant)) target specific underperforming neurons (Xu2023DrMMV) and provide a finer-grained approach by selectively reinitializing a subset of neurons identified as dormant, which mitigates forgetting and maintains computational efficiency.
@@
The effectiveness of neuron-dependent resets hinges critically on having a reliable criterion to identify which neurons require initialization. Existing methods primarily rely on activation-based metrics, such as the $\tau$ dormant neuron metric (sokar2023dormant), which measures neuronal inactivity based on activation values (i.e., the output of activation functions). It has demonstrated utility in standard architectures, and has been used to guide targeted neuron resets to restore activity in simple settings such as serial MLPs with ReLU activations (Agarap2018DeepLU), providing a simple means of maintaining learning capability without relying on auxiliary networks (nikishin2023deep) or models (Lee2024SlowAS). By identifying neurons with weak or no activation and selectively reinitializing them, the activation-based dormancy metrics become a widely adopted tool for restoring learning capabilities and preventing performance plateau in standard deep RL settings (farias2025selfnormalized; liu2025neuroplastic; juliani2024a).
@@
Figure 1: Dormant neuron metric struggles in advanced vision RL BRO-net agent (nauman2024bigger). Neuron resetting based on the dormant neuron index (ReDo) cannot restore the learning capacity of the agent, which limits its effectiveness. The curves record normalized score across 3 image-input tasks (15 runs per method), i.e., Dog Stand, Dog Walk, Dog Run.
@@
However, deep RL architectures have rapidly evolved beyond these simple architectures, integrating advanced components such as residual connections (nauman2024bigger; lee2025simba), mixture of experts (MoEs) (obando2024mixtures; sokar2024don; willimixture), and diffusion-based models (ren2025diffusion; wang2024diffusion) to improve scalability and performance for tackling complex tasks involving continuous control and high-dimensional observations. Additionally, they also imperceptibly alter the statistical properties of neuron activations, creating a significant mismatch with traditional activity metrics. Through our extensive analysis, we find that resetting based on $\tau$-dormant neuron ratio (ReDo) reduces effectiveness in these newer architectures.
As shown in Fig. 1, ReDo struggles to identify inactive neurons that have lost their learning capacity in the more advanced BRO-net architecture, which features residual paths and layer normalization, ultimately failing to achieve the intended enhancement. This mismatch arises because activation-based metrics focus solely on a neuron’s current output, evaluating its expressive capacity (how strongly it activates), while neglecting its learning capacity (how effectively it can adapt to new data distributions). Consequently, it becomes particularly problematic as deep RL architectures evolve beyond simple feedforward networks to incorporate advanced structural elements, diverse activation functions, and sophisticated normalization techniques. In these modern architectures, a neuron’s activation magnitude often fails to accurately reflect its true learning potential.
@@
To address this mismatch, we propose a fundamental shift in perspective: from evaluating neurons based on their outputs to evaluating their learning potential by leveraging gradient magnitude. While activation values only capture what a neuron currently expresses, gradients tend to measure a neuron’s capacity in response to the given situation that directly drives parameter updates. This makes it a general and natural proxy for neuronal health across architectural variations. Building upon this insight, we introduce GraMa (Gradient Magnitude based Neuronal Activity Metric), a robust and lightweight framework for quantifying neuronal activity via the gradient magnitudes. GraMa maintains validity across diverse architectural patterns, making it well-suited for modern deep RL agents.
In addition, GraMa imposes negligible computational and memory overhead by utilizing information already present in the optimization pipeline, which is a critical consideration for resource-intensive RL training. Leveraging GraMa’s efficiency, we develop a targeted neuron reset mechanism, (ReGraMa), that selectively reinitializes inactive neurons that have lost their learning capacity during training. This mechanism demonstrates robust efficiency across various architectures, including the SAC variant (nadimpalli2025evolving), the residual BRO-net (nauman2024bigger), and the diffusion-based policy DACER (wang2024diffusion).
Our contributions are summarized as follows: We show that the widely-adopted activation-based neuronal health measurements lose statistical power in complex architectures and provide a qualitative analysis of the underlying causes. We reframe neuronal health evaluation through Grama, a gradient-based metric that quantifies learning potential independently of architectural complexity, and demostrate that neuronal learning capacity degradation affects even state-of-the-art network developments. We develop (ReGraMa), an efficient neuron resetting mechanism guided by GraMa, which effectively restores neuronal activity across a wide range of network architectures. We conduct extensive experiments on MuJoCo (Brockman2016OpenAIG), DeepMind Control Suite (Tassa2018DeepMindCS), showing that GraMa-guided resetting improves performance and learning stability across diverse architectures.
@@ 2 Background
[…]
@@
Recent studies have identified that the dynamic and non-stationary nature of RL objectives can cause neurons to permanently lose their activity (Lu2018CollapseOD), impairing the network’s ability to fit new data and thereby limiting learning progress (Nikishin2022ThePB; ceron2024value). This phenomenon, referred to as the dormant neuron phenomenon, has been quantitatively characterized using the $\tau$-dormant neuron ratio (sokar2023dormant), and serves as a core metric for assessing neuron-level plasticity (Xu2023DrMMV; qin2024the; liu2025neuroplastic). Specifically, a neuron $i$ in layer $\ell$ is considered $\tau$-dormant if its normalized activation (see Eq. 1) falls below a threshold $\tau$, where $H^{\ell}$ is the number of neurons in layer $\ell$, and $D$ denotes the data distribution. $h_{i}^{\ell}(x)$ represents the activation value of neuron $i$ given input $x$.
@@
3 Related Work
Dynamic objectives may cause irreversible damage to the neuronal activity of networks during training (understand; overfit; castanyer2025stablegradientsstablelearning), which is also significant in the field of multi-agent RL (qin2024the) and visual deep RL (ma2024revisiting). This issue may lead to the phenomenon where deep learning agents progressively lose their ability to fit new data (Abbas2023LossOP), and limit their capacity for continual learning (elsayed2024addressing). In the field of Deep RL, there are some factors that have been found to have a direct correlation with the activity of neurons (juliani2024a; mayor2025the; castanyer2025stablegradientsstablelearning). Among these, activation functions play a particularly important role, recent findings indicate that replacing ReLU activation can help preserve neuron-level learning capacity (Abbas2023LossOP).
Certain loss functions, especially those incorporating L2 regularization, have been found to mitigate the network’s activity degradation (Kumar2023MaintainingPI). Another relevant line of research explores architectural adjustments such as scaling up (nikishin2023deep) or topology growth (liu2025neuroplastic). These methods can introduce modules that enhance the proportion of active, high-quality neurons. A separate and increasingly influential line of research focuses on directly manipulating model parameters to recover their activity. Resetting neural networks, either periodically or selectively, has been shown to be both effective and easy to implement (farias2025selfnormalized). For instance, Nikishin2022ThePB demonstrates the benefits of periodic resets for continual learning.
Similarly, ma2024revisiting shows that selectively resetting specific layers can improve stability without sacrificing expressive capacity. At a finer granularity, ReDO (sokar2023dormant) introduces a neuron-level resetting scheme that outperforms earlier coarse strategies in terms of stability and precision. However, ReDO’s reliance on activation-based quantization limits its applicability to more complex architectures (understand; Lyle2024DisentanglingTC). To address this limitation, we introduce a novel neuronal activity quantification approach based on gradient magnitudes, enabling generalized activity estimation and recovery across arbitrary network designs. Notably, Ji2024ACEO provided the first empirical evidence in model-free reinforcement learning of a strong correlation between the internal gradient dynamics of the agent’s policy network and its learning capacity, offering compelling motivation for the present study.
@@ 4 Gradient Magnitude based Neuronal Activity Metric (GraMa)
In this section, we qualitatively investigate ReDo’s limitations from the perspective of architectural composition Sec. 4.1. We then introduce GraMa (Sec. 4.2), a novel neuronal activity metric that redefines the statistical objective from activation values to gradient magnitudes. Next, we perform an in-depth analysis of the characteristics of low learning capacity neurons using GraMa, and provide both theoretical and empirical analyses of the performance similarity between GraMa and ReDo on simple architectures. Finally, we analyze the advantages of GraMa in complex architectures.
@@ 4.1 Misalignment of activation-based neuron activity detection in modern architectures
The widely adopted activation-based neuron activity measures (sokar2023dormant) operate on the assumption that a neuron’s activation magnitude directly corresponds to its contribution to learning. Motivated by the development of more complex network structures in language and vision domains, recent trends in deep RL have moved beyond simple serial MLP architectures with ReLU activations towards more sophisticated network designs for scaling. However, this assumption becomes increasingly fragile in modern complex architectures which are rapidly becoming standard practice. We reveal a misalignment between activation values and actual learning potential, leading to inefficient neuron identification.
@@
To systematically analyze this misalignment, we evaluate BRO-net (nauman2024bigger), a recently proposed architecture, on the challenging Dog Walk task (Tassa2018DeepMindCS). We categorized neurons into four quadrants based on their expressive capacity (measured by activation magnitude) and learning capacity (measured by gradient magnitude), which we introduce in more detail in Sec. 4.2. Neurons are ranked according to these two criteria, and we select the top 25% of neurons in each ranking, so as to provide an understanding of the types of neurons reset by each method. The results in Fig. 3 demonstrate that ReDo cannot accurately identify neurons that have genuinely lost their learning capacity (red) based on their activation values. More importantly, ReDo mistakenly resets a significant number of neurons with high learning capacity that are less expressive at the moment (dark blue).
@@
In addition to the primary factor of statistical objectives, we identify three key architectural features that may undermine the reliability of activation-based neuron dormancy detection. Fig. 2 provides an intuitive illustration of each. (i) Multi-branch network structures. In modern architectures like ResNets (He2015DeepRL), information flows through multiple branches before being fused and passing through activation functions. This architectural pattern introduces a critical problem: the final activation values observed after branch fusion no longer accurately reflect the individual contributions of neurons in each branch. Our experiments with Resnet-SAC (Shah2021RRLRA) shown in Fig. 4 (a) confirm this effect, where the results suggest that activation-based methods lose their ability to identify dormant neurons and limits the performance of the agent. (ii) Non-ReLU activation function. The interpretability of activation values is highly dependent on the specific activation function used. While ReLU creates a clear distinction between “dead” (output=0) and “active” neurons, this clarity breaks down when considering alternative activation functions. For example, with Leaky ReLU, neurons may remain expressive and contribute to learning even when their pre-activation values are negative.
As a result, solely relying on activation values to measure neuron dormancy becomes unreliable and may lead to misidentification. Our experiments replacing ReLU with Leaky ReLU in SAC (Fig. 4 (b)) demonstrate this problem, where the activation-based method (ReDo) shows considerable reset activity during early learning with lower performance, indicating that it struggles to establish meaningful dormancy criteria. (iii) Normalization layers. Normalization techniques modify the distribution of neuron outputs before they pass through activation functions. This process adjusts outliers and rescales values across entire layers, causing post-normalization values to lose a direct correspondence with individual neuron functionality. Thus, activation values measured after normalization may no longer accurately reflect a neuron’s learning capacity. Our experiments incorporating layer normalization into SAC, Fig. 4 (c) reveal that the activation-based method struggles to maintain consistent neuron assessment, whereas our method consistently identifies relevant neurons throughout the learning process.
Figure 4: Empirical validation corresponding to the three cases of Fig. 2. Top row records the proportion of reset neurons. Bottom row shows the performance. Each curve represents the average over 3 seeds (Dog Walk).